An arrow is fired directly upwards at 15m/s from gorund level.
A) How long is the arrow in the air?
B) How high does the arrow go?
C) How fast is it going just before it strikes the ground?
1
Expert's answer
2018-12-05T15:29:10-0500

Velocity:

V=15m/s

Gravity:

g=9.8 m/s^2


Solution:

Vertical velocity at maximum height is 0:

V-gt=0

Hence the time to reach the maximum height:

t=V/g=15/9.8=1.53s

According to the law of motion, the maximum height is:

h=Vt-(gt^2)/2=15*1.53-9.8*〖1.53〗^2/2=11.5m

Since the time of fall is equal to the time of take-off, the total time the arrow stays in the air until it touches the ground is:

T=2t=1.53*2=3.1s